OpenStudy (anonymous):

whats the sum of this infinite series: 9/2 - 3/4 + 1/8 - 1/48 + 1/288...

hartnn (hartnn):

i got 27/7 its the difference of sum of two series {9/2+1/8+...}and{3/4+1/48+....}

OpenStudy (anonymous):

the formula is initial value * 1/(1-ratio between precedent and next number)

OpenStudy (anonymous):

which is 9/2 * 1/(1- -1/8) right?

hartnn (hartnn):

yup or in simple terms a/(1-r) a is 1st term,r is common ratio... i got r=1/36 for both series. how 1/8 ??

OpenStudy (anonymous):

its 1/6?

hartnn (hartnn):

ratio=(1/8)/(9/2)=1/36

OpenStudy (anonymous):

THANKS U ARE RIGHT

hartnn (hartnn):

ratio for 2nd series=(1/48)/(3/4)=1/36 so u got final answer?

OpenStudy (anonymous):

YES U ARE A GENIUS

hartnn (hartnn):

nopes,i just practice math :)